Find an equation of the line that has a y intercept of -2 and is perpendicular to the graph of 3x + 6y=2
Answer by Alan3354(69443)

Find an equation of the line that has a y intercept of -2 and is perpendicular to the graph of 3x + 6y=2
---------------------
Ok. First, find the slope (m) of 3x+6y = 2 by putting it into slope-intercept form (y = mx + b).
3x+6y = 2
6y = -3x + 2
y = (-1/2)x + 1/3
The slope is -1/2. Lines perpendicular will have the negative inverse slope, or +2.
Now we know the slope (2) and the y-intercept is -2, so that's the point (0,-2).
y-y1 = m*(x-x1) where (x1,y1) is the point (0,-2)
y+2 = 2*(x-0)
y+2 = 2x
Or, 2x-y = 2
In slope intercept form, it's y = 2x - 2, so the y-intercept is 2, as spec'd.